how do you solve this system of equation
x+y=8
51x+8y=193

Answer by nerdybill(7384)

You can solve by one of three methods:
Substitution, addition or matrix.
.
I'll apply the "addition method":
x+ y= 8
51x+8y=193
.
Multiply first equation by -8 to get:
-8x-8y=-64
51x+8y=193
.
Add the two equations together:
-8x-8y=-64
51x+8y=193
-----------
43x = 129
x = 129/43
x = 3
.
To find y, substitute the above into:
x+ y= 8
3+ y= 8
y = 5
.
Our solution:
x = 3
y = 5